OpenStudy (anonymous):

Please help! The bread machine you are interested in costs $100 with tax. The ingredients to make one loaf of bread cost $0.80. What is the rate of cost of (one) loaf of bread? y=mx+b

OpenStudy (anonymous):

hum... total cost = 100+0.8n, div that by n, cost per bread = (100+0.8n)/n=100/n +0.8

OpenStudy (anonymous):

btw n is the number of breads